# Clustering of timed sequences – Application to the analysis of care pathways

### v2

@@ 7 Application to Real-World Care Pathways
In this section, we apply the proposed method to real-world care pathways and compare our results with those of Sequence State Analysis (SSA) (Gabadinho et al., 2011). SSA, particularly the TraMineR tool, is commonly used for categorizing care pathways (Roux et al., 2019; Le Meur et al., 2019).
@@
For this experiment, we used the servers from the Paris University Hospital made of AMD EPYC 7713 processors. Our algorithms are neither parallelized nor distributed. However, the implementation of softDTW by Tavenard et al. (2020) can utilize multiple cores. During our experiments, we requested only one core from the servers.
@@ 7.1 Comparison with TraMineR
Figure 3 corresponds to the $K=5$ clusters for the three methods. Each cluster is represented as an histogram of events over time. One can first notice that TraMineR provides clusters that look very similar to each others. It is difficult to identify specificities for each cluster. Contrary to the other two methods, based on our the drop-DTW metric, the sizes of the clusters are more homogeneous. For HierAsTiSeq and K-means, there is one large cluster and several small clusters that gather patients specific timed events. The two large clusters (cluster 1 for HierAsTiSeq and cluster 1 for K-means) looks very similar in distribution of events. HierAsTiSeq identifies two clusters with specific behaviors: the cluster 3 has no post-operative events. This may correspond to patients that had the surgery in the hospital but their follow-up in another medical structure; the cluster 2 are patients with short delays between pre- or post-operative care events around the surgery. The clusters 3 and 4 identified by K-means also gather patients with a late pre-operative consultation (wrt the resection). According to the distributions, these patients have two pre-operative consultations: in cluster 3, the first pre-operative consultation occurs earlier than in cluster 4; and cluster 4 holds patients having a late post-operative consultation.
We conclude from this experiment that the use of drop-DTW yields clusters of patients with more specific timed sequences than TraMineR. For this dataset, the latter method does not seems to be able to reveal interesting clusters of patients. We performed a quantitative (confusion matrices) and qualitative analysis. The analysis of confusions between HierAsTiSeq and TraMineR shows that there is no correspondence between the clusters. This suggests that the metrics are semantically different. In this case, the semantics of drop-DTW on timed sequence seems more meaningful, but it does not mean that it would be the case for all datasets.
It is worth noting that results of K-means with drop-DTW may significantly differ across different runs due to the randomness of the cluster seeds. HierAsTiSeq and TraMineR are more deterministic, but the iterative computation of drop-DTW or optimal matching may introduce some randomness. Nonetheless, the hierarchical clustering strategy of HierAsTiSeq and TraMineR is resource demanding. This prevents to apply them on the entire dataset. This was possible only with K-means and takes only a few minutes.
@@ 7.2 Cluster Analysis of K-means with drop-DTW
We can also notice that there is a clear alignment between the barycenters events and the event distribution. This illustrates that the averaging method, based on drop-DTW provides a meaningful average of a collection of complex timed sequences. More specifically, comparing the clusters by the decay between consultations and surgery leads to the same conclusions as before.
@@ 8 Conclusion
We proposed a method for clustering timed sequences inspired by existing techniques for time series. Thus, the use of timed sequences allows co-occurring events, irregular time sampling, and sparse events. All these advantages are very interesting from a clinical point of view for describing care pathways. We showed that by using drop-DTW, the clustering algorithm are less sensitive to missing and aberrant values when drop-DTW is used. The result is interesting for application to real-world care pathways. Then, we tested our method on real data from the OPTISOINS study for our approach compared to TraMineR. Based on the initial results, drop-DTW seems to be clinically promising for identifying average pathways.

### v3

@@ 7 Application to Real-World Care Pathways
In this section, we apply the proposed method to real-world care pathways and compare our results with those of Sequence State Analysis (SSA) [8]. SSA, particularly the TraMineR tool, is commonly used for categorizing care pathways [19, 14].
@@
For this experiment, we used the servers from the Paris University Hospital made of AMD EPYC 7713 processors. Our algorithms are neither parallelized nor distributed. However, the implementation of softDTW by [23] can utilize multiple cores. During our experiments, we requested only one core from the servers.
@@ 7.1 Comparison with TraMineR
Figure 3 corresponds to the $K=5$ clusters for the three methods. Each cluster is represented as an histogram of events over time. One can first notice that TraMineR provides clusters that look very similar to each others. It is difficult to identify specificities for each cluster. Contrary to the other two methods, based on our drop-DTW metric, the sizes of the clusters are more homogeneous. For HierAsTiSeq and K-means, there is one large cluster and several small clusters that gather patient specific timed events. The two large clusters (cluster 1 for HierAsTiSeq and cluster 1 for K-means) looks very similar in distribution of events. HierAsTiSeq identifies two clusters with specific behaviors: the cluster 3 has no post-operative events. This may correspond to patients that had a surgery in the hospital but their follow-up in another medical structure; the cluster 2 are patients with short delays between pre- or post-operative care events around the surgery. The clusters 3 and 4 identified by K-means also gather patients with a late pre-operative consultation (wrt the resection). According to the distributions, these patients have two pre-operative consultations: in cluster 3, the first pre-operative consultation occurs earlier than in cluster 4; and cluster 4 holds patients having a late post-operative consultation.
We conclude from this experiment that the use of drop-DTW yields clusters of patients with more specific timed sequences than TraMineR. For this dataset, the latter method does not seem to be able to reveal interesting clusters of patients. We performed a quantitative (confusion matrices) and qualitative analysis. The analysis of confusion between HierAsTiSeq and TraMineR shows that there is no correspondence between the clusters. This suggests that the metrics are semantically different. In this case, the semantics of drop-DTW on timed sequence seems more meaningful, but it does not mean that it would be the case for all datasets.
It is worth noting that results of K-means with drop-DTW may significantly differ across different runs due to the randomness of the cluster seeds. HierAsTiSeq and TraMineR are more deterministic, but the iterative computation of drop-DTW or optimal matching may introduce some randomness. Nonetheless, the hierarchical clustering strategy of HierAsTiSeq and TraMineR is resource demanding. This prevents applying them on the entire dataset. This was possible only with K-means and takes only a few minutes.
@@ 7.2 Cluster Analysis of K-means with drop-DTW
We can also notice that there is a clear alignment between the barycenters’ events and the event distribution. This illustrates that the averaging method, based on drop-DTW provides a meaningful average of a collection of complex timed sequences. More specifically, comparing the clusters through the decay between consultations and surgery leads to the same conclusions as before.
@@ 8 Conclusion
We proposed a method for clustering timed sequences inspired by existing techniques for time series. Thus, the use of timed sequences allows co-occurring events, irregular time sampling, and sparse events. All these advantages are very interesting from a clinical point of view for describing care pathways. We showed that by using drop-DTW, the clustering algorithm is less sensitive to missing and aberrant values when drop-DTW is used. The result is interesting for application to real-world care pathways. Then, we tested our method on real data from the OPTISOINS study for our approach compared to TraMineR. Based on the initial results, drop-DTW seems to be clinically promising for identifying average pathways.